# Do Sparse Autoencoders Generalize? A Case Study of Answerability

### v1

@@ Acknowledgements
The authors also thank the developers of TransformerLens Nanda and Bloom (2022) and SAELens Bloom et al. (2024), open-source libraries for mechanistic interpretability.
@@ Appendix A Preliminaries
To train SAE probes with more than 1 feature, we follow the general methodology of our 1-sparse probes. As testing all possible SAE feature combinations is computationally infeasible, we iteratively increase the number of features while testing only the most promising candidates for higher features combinations. Specifically, to find combinations of $k$ features, we use the top 50 best performing features of size $k-1$ and test all possible new combinations with the 500 best performing single SAE features. We use a constant regularization parameter of 1 for the probes, regardless of the number of features.
@@
Appendix B Datasets
Full Dataset details.
  • Celebrity Recognition: Queries requiring knowledge about celebrities. For construction, we use a public dataset of actors and movies from IMDB [^2] https://www.kaggle.com/datasets/darinhawley/imdb-films-by-actor-for-10k-actors, and generate a list of the 1000 most popular actors after 1990, as measured by the total number of ratings their movies received. We construct an additional dataset of non-celebrity names by randomly generating first and last name combinations using the most common North American names from Wikipedia [^3] https://en.wikipedia.org/wiki/Lists_of_most_common_surnames_in_North_American_countries and https://en.wikipedia.org/wiki/List_of_most_popular_given_names?utm_source=chatgpt.com.
Dataset sizes
Dataset Size
SQUAD (train) 2000
BoolQ (train) 2000
SQUAD (test) 1800
SQUAD (variations) 1800
BoolQ (test) 2000
IDK 484
Equation 2000
Celebrity 600
Table 2: Number of examples for each used dataset.
Table 2 shows the number of examples for each dataset used in our evaluation.
Appendix C Additional analysis
C.1 Answerability Detection at Different Layers
@@
We repeat our SAE feature analysis in Layer 20 of the model, as well as providing additional analysis for SAE features activations sampled after the activation function. Figure 6 shows the Layer 20 results using activations sampled before the activation function, while Figures 5 and 7 show analogous results when sampling SAE activations after the activation function. Sampling after the activation reduces the number of relevant features our probe finds, since many features are inactive. However, this does not change the overall results, as we still find features with good generalization performance.
@@ C.2 Prompt variations
C.3 In-domain SAE feature accuracies
@@
C.4 SAE Feature Combination Analyses
@@ C.5 Cosine Similarities
Other experiments.

### v2

@@ Acknowledgements
The authors also thank the developers of TransformerLens Nanda & Bloom (2022) and SAELens Bloom et al. (2024), open-source libraries for mechanistic interpretability.
@@ Appendix A Preliminaries
To train SAE probes with more than 1 feature, we follow the general methodology of our 1-sparse probes. As testing all possible SAE feature combinations is computationally infeasible, we iteratively increase the number of features while testing only the most promising candidates for higher features combinations. Specifically, to find combinations of $k$ features, we use the top 50 best performing features of size k1k-1 and test all possible new combinations with the 500 best performing single SAE features. We use a constant regularization parameter of 1 for the probes, regardless of the number of features.
@@
Appendix B Additional analysis
B.1 Answerability Detection at Different Layers
@@
We repeat our SAE feature analysis in Layer 20 of the model, as well as providing additional analysis for SAE features activations sampled after the activation function. Figure 7 shows the Layer 20 results using activations sampled before the activation function, while Figures 6 and 8 show analogous results when sampling SAE activations after the activation function. Sampling after the activation reduces the number of relevant features our probe finds, since many features are inactive. However, this does not change the overall results, as we still find features with good generalization performance.
@@ B.2 Prompt variations
B.3 In-domain SAE feature accuracies
@@
B.4 SAE Feature Combination Analyses
@@ B.5 Cosine Similarities
B.6 Spanish language generalization
We performed an additional generalization experiment using a Spanish language variant of SQUAD Carrino et al. (2019). Similar to our other OOD experiments, we used residual and SAE probes trained on the English SQUAD dataset and evaluated their generalization capabilities to the Spanish SQUAD data. Our results did not show large differences in generalization between individual features or the residual stream probe (see Figure 15). Instead, all top probes and the residual probe generalized roughly equally well, with a performance gap of slightly below 10%.
Figure 15: Probing accuracies for top SAE features and the residual stream probe when evaluated on Spanish OOD data.
Other experiments